OpenStudy (anonymous):

Need help asap please! Write the expression as a binomial x^2-4/9x+4/81

OpenStudy (mathstudent55):

You mean this expression? \(x^2 - \dfrac{4}{9}x + \dfrac{4}{81} \)

OpenStudy (anonymous):

yes :) @mathstudent55 well the x is next to only the 9

OpenStudy (mathstudent55):

Notice that x^2 is the square of x. Also, 4/81 is the square of 2/9. This could well be the square of a binomial.

OpenStudy (mathstudent55):

If it is the square of the binomial, then you'd have and x from the x^2 and a 2/9 from the 4/81. Since the sign of the x term is negative, you'd have a minus sign. \((x - \dfrac{2}{9} )^2 \) All you need to do is square the binomial and see if the middle term of your given trinomial fits in.

OpenStudy (mathstudent55):

When you square a binomial you do this: \( (a - b)^2 = a^2 - 2ab + b^2\) We already know that the x^2 term and the number at the end fit in. Let's check the middle term. \(-2ab = -2(x)(\dfrac{2}{9}) = -\dfrac{4}{9}x\) The middle term checks out, so you can conclude that \(x^2 - \dfrac{4}{9}x + \dfrac{4}{81} = (x - \dfrac{2}{9} )^2\)

OpenStudy (anonymous):

Thank you so much! :) @mathstudent55

OpenStudy (mathstudent55):

You're welcome.